# Strong and weak $CP$ tests in sequential decays of polarized $\Sigma^{0}$ hyperons

### v1

@@ Abstract
The $J/\psi,\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ processes and subsequent decays are studied using the world’s largest $J/\psi$ and $\psi(3686)$ data samples collected with the BESIII detector. The strong-$CP$ symmetry is tested in the decays of the $\Sigma^{0}$ hyperons for the first time by measuring the decay parameters, $\alpha_{\Sigma^{0}}=-0.0017\pm 0.0021\pm 0.0018$ and $\bar{\alpha}_{\Sigma^{0}}=0.0021\pm 0.0020\pm 0.0022$. The weak-$CP$ test is performed in the subsequent decays of their daughter particles $\Lambda$ and $\bar{\Lambda}$. Also for the first time, the transverse polarizations of the $\Sigma^{0}$ hyperons in $J/\psi$ and $\psi(3686)$ decays are observed with opposite directions, and the ratios between the S-wave and D-wave contributions of the $J/\psi,\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ decays are obtained. These results are crucial to understand the decay dynamics of the charmonium states and the production mechanism of the $\Sigma^{0}-\bar{\Sigma}^{0}$ pairs.
@@
The question of why our universe consists of matter, but almost no antimatter, has puzzled the scientific community for more than half a century and is to this day subjected to intensive research Ablikim et al. (2022a, b, 2023a); Aaij et al. (2024); Guan et al. ; Aaij et al. . One of the long-standing explanations is baryogenesis Sakharov (1967), i.e. that the matter abundance is generated dynamically. Among the necessary criteria for this to be possible is the existence of processes that violate charge-conjugation and parity ($CP$) conservation. The Standard Model (SM) of particle physics allows for tiny signals of $CP$ violation ($CPV$) in weak interactions, in line with experimental observations in meson decays Christenson et al. (1964); Aubert et al. (2001); Abe et al. (2001); Aaij et al. (2019). The SM should in principle also allow for $CP$ violating strong processes, which would be manifest in e.g. a non-zero neutron electric dipole moment (EDM) Ramsey (1982); Fortson et al. (2003); Luders (1957). However, the extremely tiny experimental upper limit of the neutron EDM Abel et al. (2020) suggests that $CPV$ in strong interactions is unnaturally small. The SM itself does not provide an answer to the smallness of strong $CPV$, which calls for further investigations. The radiative decay $\Sigma^{0}\to\Lambda\gamma$ could in principle offer an opportunity to study the interference between a parity-conserving (from the magnetic transition moment) and a parity-violating amplitude (from the electric dipole transition moment, EDTM). The EDTM of this decay is related to the neutron EDM via SU(3) flavor symmetry Gell-Mann (1961); Nair et al. (2019). This can be explored by measuring the electromagnetic decay parameters $\alpha_{\Sigma^{0}}$, and $\bar{\alpha}_{\Sigma^{0}}$ for the charge conjugate decay; any non-zero value of $\alpha_{\Sigma^{0}}/\bar{\alpha}_{\Sigma^{0}}$ would point to signals of $CPV$ beyond the SM Nair et al. (2019).
@@
An unbinned maximum likelihood fit is performed in the seven angular dimensions $\vec{\zeta}$ by simultaneously fitting both the $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ datasets to determine the parameters $\vec{\omega}$. The numerical results are summarized in Table 1, together with the $CP$ asymmetry in the $\Lambda$ decay, $A^{\Lambda}_{CP}=(\alpha_{\Lambda}+\bar{\alpha}_{\Lambda})/(\alpha_{\Lambda}-\bar{\alpha}_{\Lambda})$.
Table 1: Values and uncertainties of the fit parameters and the $CP$ asymmetry $A^{\Lambda}_{CP}$, along with the previous measurements. The first and second uncertainties in this work are statistical and systematic, respectively, and those for the previous results are the total uncertainties. In the previous results, only the values reported in Ref. Ablikim et al. (2017) are cited for $\alpha_{J/\psi}$ and $\alpha_{\psi(3686)}$, as it gives the most precise measurement for $\alpha_{J/\psi}$ and the only measurement for $\alpha_{\psi(3686)}$.
@@
where $n_{1,x}=\cos\theta_{p}\sin\theta_{\Lambda}\cos\varphi_{\Lambda}$, $n_{1,y}=\cos\theta_{p}\sin\theta_{\Lambda}\sin\varphi_{\Lambda}$, $n_{1,z}=\cos\theta_{p}\cos\theta_{\Lambda}$, $n_{2,x}=\cos\theta_{\bar{p}}\sin\theta_{\bar{\Lambda}}\cos\varphi_{\bar{\Lambda}}$, $n_{2,y}=\cos\theta_{\bar{p}}\sin\theta_{\bar{\Lambda}}\sin\varphi_{\bar{\Lambda}}$, $n_{2,z}=\cos\theta_{\bar{p}}\cos\theta_{\bar{\Lambda}}$. Figure 3 shows the distributions of the moments versus $\cos\theta_{\Sigma}$ for the $J/\psi$ dataset. The fit results are consistent with those distirbutions in data.
@@
Source ($10^{-3}$) $\alpha_{J/\psi}$ $\Delta\Phi_{J/\psi}$ (rad) $\alpha_{\psi(3686)}$ $\Delta\Phi_{\psi(3686)}$ (rad) $\alpha_{\Sigma^{0}}$ $\bar{\alpha}_{\Sigma^{0}}$ $\alpha_{\Lambda}$ $\bar{\alpha}_{\Lambda}$ $A^{\Lambda}_{CP}$
Tracking 0.0 0.2 10 13 0.9 1.5 1 1 0
$\gamma$ reconstruction 0.5 0.5 1 0 0.2 0.1 2 3 3
Λ/$\Lambda/\bar{\Lambda}$ reconstruction 6.2 0.2 24 30 0.7 0.4 6 3 8
4C kinematic fit 1.5 0.1 5 6 1.0 1.5 2 3 3
$\Sigma^{0}/\bar{\Sigma}^{0}$ mass window 0.2 3.2 2 7 0.8 0.2 7 6 8
Background estimation 4.3 0.4 10 6 0.6 0.1 6 6 9
Total 7.7 3.3 28 34 1.8 2.2 11 10 15
@@
In summary, this Letter presents the first strong-$CP$ tests in $\Sigma^{0}\to\Lambda\gamma$ and $\bar{\Sigma}^{0}\to\bar{\Lambda}\gamma$ decays by measuring the decay parameters $\alpha_{\Sigma^{0}}=-0.0017\pm 0.0021\pm 0.0018$ and $\bar{\alpha}_{\Sigma^{0}}=0.0021\pm 0.0020\pm 0.0022$, respectively. These results are consistent with $CP$ conservation, and will provide crucial constraints for new physics models. The decay parameters $\alpha_{\Lambda}$ and $\bar{\alpha}_{\Lambda}$, which are listed in Table 1, are measured independently, and the weak-$CP$ test is performed in the sub-decays of $\Lambda$ and $\bar{\Lambda}$. These results are in good agreement with the values obtained from the $J/\psi\to\Lambda\bar{\Lambda}$ and $J/\psi\to\Xi^{-(0)}\bar{\Xi}^{+(0)}$ analyses of BESIII Ablikim et al. (2022a, b, 2023a).
Furthermore, the spin polarizations of the $\Sigma^{0}$ hyperons with opposite directions in the $J/\psi$ and $\psi(3686)$ decays are observed for the first time. This phenomenon is also observed in the case of the $\Sigma^{+}$ hyperon decays Ablikim et al. (2020a), but not in the $\Xi^{-(0)}$ Ablikim et al. (2023a, b, 2022a, 2022c) hyperon decays. Following Ref. Wu et al. (2021) with the measured parameters $\alpha_{\psi}$ and $\Delta\Phi_{\psi}$, the ratio of the D-wave and S-wave coupling constants $g_{D}/g_{S}$, the relative phase $\delta$ between the S-wave and D-wave, the percentage of the S-wave contribution $\Gamma_{S}/\Gamma_{\rm total}$, and the effective radius $r_{\rm eff}$ of the $\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ decay (where $\bar{L}=r_{\rm eff}\times p$, $\bar{L}$ and $p$ are the average orbital angular momentum and the relative momentum of the $\Sigma^{0}$ and $\bar{\Sigma}^{0}$ in $\psi$ C.M. system) are obtained for the first time, as shown in Table 3 together with the parameters for $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ calculated in Ref. Wu et al. (2021). What is particularly intriguing is that, as speculated in Ref. Wu et al. (2021), the $\delta$ difference between $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ is about the same as the value between $J/\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ and $\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ decays Wu et al. (2021), and both are equal to approximately $\pi$. If this is not a coincidence, the underlying mechanism should be investigated in the future. These results are crucial to understand the decay dynamics of the charmonium states and the production mechanism of the $\Sigma^{0}-\bar{\Sigma}^{0}$ pairs and provide useful information for hunting down excited nucleon resonances Wu et al. (2021).
@@ Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. The authors would like to extend thanks to Dr. Shu-Ming Wu and Dr. Jia-Jun Wu for useful discussion and helpful advice. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406300, 2020YFA0406400; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12165022, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12342502, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; Yunnan Fundamental Research Project under Contract No. 202301AT070162; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; The Knut and Alice Wallenberg Foundation, Sweden; The Swedish Foundation for International Cooperation in Research and Higher Education (STINT); U. S. Department of Energy under Contract No. DE-FG02-05ER41374.

### v2

@@ Abstract
The $J/\psi,\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ processes and subsequent decays are studied using the world’s largest $J/\psi$ and $\psi(3686)$ data samples collected with the BESIII detector. The parity-violating decay parameters of the decays $\Sigma^{0}\to\Lambda\gamma$ and $\bar{\Sigma}^{0}\to\bar{\Lambda}\gamma$, $\alpha_{\Sigma^{0}}=-0.0017\pm 0.0021\pm 0.0018$ and $\bar{\alpha}_{\Sigma^{0}}=0.0021\pm 0.0020\pm 0.0022$, are measured for the first time. The strong-$CP$ symmetry is tested in the decays of the $\Sigma^{0}$ hyperons for the first time by measuring the asymmetry $A^{\Sigma}_{CP}=\alpha_{\Sigma^{0}}+\bar{\alpha}_{\Sigma^{0}}=(0.4\pm 2.9\pm 1.3)\times 10^{-3}$. The weak-$CP$ test is performed in the subsequent decays of their daughter particles $\Lambda$ and $\bar{\Lambda}$. Also for the first time, the transverse polarizations of the $\Sigma^{0}$ hyperons in $J/\psi$ and $\psi(3686)$ decays are observed with opposite directions, and the ratios between the S-wave and D-wave contributions of the $J/\psi,\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ decays are obtained. These results are crucial to understand the decay dynamics of the charmonium states and the production mechanism of the $\Sigma^{0}-\bar{\Sigma}^{0}$ pairs.
@@
The question of why our universe consists of matter, but almost no antimatter, has puzzled the scientific community for more than half a century and is to this day subjected to intensive research Ablikim et al. (2022a, b, 2023a); Aaij et al. (2024); Guan et al. ; Aaij et al. . One of the long-standing explanations is baryogenesis Sakharov (1967), i.e. that the matter abundance is generated dynamically. Among the necessary criteria for this to be possible is the existence of processes that violate charge-conjugation and parity ($CP$) conservation. The Standard Model (SM) of particle physics allows for tiny signals of $CP$ violation ($CPV$) in weak interactions, in line with experimental observations in meson decays Christenson et al. (1964); Aubert et al. (2001); Abe et al. (2001); Aaij et al. (2019). The SM should in principle also allow for $CP$ violating strong processes, which would be manifest in e.g. a non-zero neutron electric dipole moment (EDM) Ramsey (1982); Fortson et al. (2003); Luders (1957). However, the extremely tiny experimental upper limit of the neutron EDM Abel et al. (2020) suggests that $CPV$ in strong interactions is unnaturally small. The SM itself does not provide an answer to the smallness of strong $CPV$, which calls for further investigations. The radiative decay $\Sigma^{0}\to\Lambda\gamma$ could in principle offer an opportunity to study the interference between a parity-conserving (from the magnetic transition moment) and a parity-violating amplitude (from the electric dipole transition moment, EDTM). The EDTM of this decay is related to the neutron EDM via SU(3) flavor symmetry Gell-Mann (1961); Nair et al. (2019). This can be explored by measuring the parity-violating decay parameters $\alpha_{\Sigma^{0}}$, and $\bar{\alpha}_{\Sigma^{0}}$ for the charge conjugate decay. In the SM, $\alpha_{\Sigma^{0}}$ and $\bar{\alpha}_{\Sigma^{0}}$ are predicted to be very small Nair et al. (2019), nonzero measurements of these parameters at current or near-future experimental sensitivity levels would be indications of parity violation beyond the standard model. Based on the values of these parameters, the strong-$CP$ symmetry can be tested by measuring the asymmetry $A^{\Sigma}_{CP}=\alpha_{\Sigma^{0}}+\bar{\alpha}_{\Sigma^{0}}$, finding a non-vanishing $A^{\Sigma}_{CP}$ would point to signals of $CPV$ beyond the SM Nair et al. (2019).
@@
An unbinned maximum likelihood fit is performed in the seven angular dimensions $\vec{\zeta}$ by simultaneously fitting both the $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ datasets to determine the parameters $\vec{\omega}$. During the fitting process, the background events estimated from the data sidebands are included, and their log-likelihood is subtracted from the data based on the normalized weights. The numerical results are summarized in Table 2, together with the $CP$ asymmetries $A^{\Sigma}_{CP}$ and $A^{\Lambda}_{CP}=(\alpha_{\Lambda}+\bar{\alpha}_{\Lambda})/(\alpha_{\Lambda}-\bar{\alpha}_{\Lambda})$. The correlation coefficient matrix of the fitted parameters is shown in Tab. 1.
Table 1: Full correlation coefficient matrix of the fitted parameters.
Coefficients $\alpha_{J/\psi}$ $\Delta\Phi_{J/\psi}$ $\alpha_{\psi(3686)}$ $\Delta\Phi_{\psi(3686)}$ $\alpha_{\Sigma^{0}}$ $\bar{\alpha}_{\Sigma^{0}}$ $\alpha_{\Lambda}$ $\bar{\alpha}_{\Lambda}$
$\alpha_{J/\psi}$ 1.000 -0.016 -0.012 0.012 -0.002 -0.002 0.018 -0.013
$\Delta\Phi_{J/\psi}$0.016 1.000 0.006 -0.007 0.004 0.000 0.011 0.027
$\alpha_{\psi(3686)}$0.012 0.006 1.000 -0.451 0.007 -0.007 -0.011 0.004
$\Delta\Phi_{\psi(3686)}$ 0.012 -0.007 -0.451 1.000 -0.001 -0.001 -0.051 -0.067
$\alpha_{\Sigma^{0}}$0.002 0.004 0.007 -0.001 1.000 -0.009 -0.083 -0.083
$\bar{\alpha}_{\Sigma^{0}}$0.002 $0$0.007 -0.001 -0.009 1.000 0.032 0.033
$\alpha_{\Lambda}$ 0.018 0.011 -0.011 -0.051 -0.083 0.032 1.000 0.980
$\bar{\alpha}_{\Lambda}$0.013 0.027 0.004 -0.067 -0.083 0.033 0.980 1.000
Table 2: Values and uncertainties of the fit parameters and the $CP$ asymmetries $A^{\Sigma}_{CP}$ and $A^{\Lambda}_{CP}$, along with the previous measurements. The first and second uncertainties in this work are statistical and systematic, respectively, and those for the previous results are the total uncertainties. In the previous results, only the values reported in Ref. Ablikim et al. (2017) are cited for $\alpha_{J/\psi}$ and $\alpha_{\psi(3686)}$, as it gives the most precise measurement for $\alpha_{J/\psi}$ and the only measurement for $\alpha_{\psi(3686)}$.
@@
$A^{\Sigma}_{CP}$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}(0.4\pm 2.9\pm 1.3)\times 10^{-3}$
@@
where $n_{1,x}=\cos\theta_{p}\sin\theta_{\Lambda}\cos\varphi_{\Lambda}$, $n_{1,y}=\cos\theta_{p}\sin\theta_{\Lambda}\sin\varphi_{\Lambda}$, $n_{1,z}=\cos\theta_{p}\cos\theta_{\Lambda}$, $n_{2,x}=\cos\theta_{\bar{p}}\sin\theta_{\bar{\Lambda}}\cos\varphi_{\bar{\Lambda}}$, $n_{2,y}=\cos\theta_{\bar{p}}\sin\theta_{\bar{\Lambda}}\sin\varphi_{\bar{\Lambda}}$, $n_{2,z}=\cos\theta_{\bar{p}}\cos\theta_{\bar{\Lambda}}$. Figure 3 shows the distributions of these moments versus $\cos\theta_{\Sigma}$ for the $J/\psi$ dataset, the ones for the $\psi(3686)$ dataset are shown in Ref. 4. The fit results are consistent with those distirbutions in data.
[…]
@@
Source ($10^{-3}$) $\alpha_{J/\psi}$ $\Delta\Phi_{J/\psi}$ (rad) $\alpha_{\psi(3686)}$ $\Delta\Phi_{\psi(3686)}$ (rad) $\alpha_{\Sigma^{0}}$ $\bar{\alpha}_{\Sigma^{0}}$ $\alpha_{\Lambda}$ $\bar{\alpha}_{\Lambda}$ $A^{\Sigma}_{CP}$ $A^{\Lambda}_{CP}$
Tracking 0.0 0.2 10 13 0.9 1.5 1 1 0.6 0
$\gamma$ reconstruction 0.5 0.5 1 0 0.2 0.1 2 3 0.0 3
Λ/$\Lambda/\bar{\Lambda}$ reconstruction 6.2 0.2 24 30 0.7 0.4 6 3 0.2 8
4C kinematic fit 1.5 0.1 5 6 1.0 1.5 2 3 0.6 3
$\Sigma^{0}/\bar{\Sigma}^{0}$ mass window 0.2 3.2 2 7 0.8 0.2 7 6 0.8 8
Background estimation 4.3 0.4 10 6 0.6 0.1 6 6 0.6 9
Total 7.7 3.3 28 34 1.8 2.2 11 10 1.3 15
@@
In summary, this Letter presents the first measurements of the parity-violating decay parameters for the decays $\Sigma^{0}\to\Lambda\gamma$ and $\bar{\Sigma}^{0}\to\bar{\Lambda}\gamma$, $\alpha_{\Sigma^{0}}=-0.0017\pm 0.0021\pm 0.0018$ and $\bar{\alpha}_{\Sigma^{0}}=0.0021\pm 0.0020\pm 0.0022$, respectively. The extracted values of decay parameters allow for the strong-$CP$ test, $A^{\Sigma}_{CP}=(0.4\pm 2.9\pm 1.3)\times 10^{-3}$. These results are consistent with $P$ and $CP$ conservation, and will provide crucial constraints for new physics models. The decay parameters $\alpha_{\Lambda}$ and $\bar{\alpha}_{\Lambda}$, which are listed in Table 2, are measured independently, and the weak-$CP$ test is performed in the sub-decays of $\Lambda$ and $\bar{\Lambda}$. These results are in good agreement with the values obtained from the $J/\psi\to\Lambda\bar{\Lambda}$ and $J/\psi\to\Xi^{-(0)}\bar{\Xi}^{+(0)}$ analyses of BESIII Ablikim et al. (2022a, b, 2023a).
Furthermore, the spin polarizations of the $\Sigma^{0}$ hyperons with opposite directions in the $J/\psi$ and $\psi(3686)$ decays are observed for the first time. This phenomenon is also observed in the case of the $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ decays Ablikim et al. (2020a), but not in the $\Xi^{-(0)}$ Ablikim et al. (2023a, b, 2022a, 2022c) hyperon pairs. Following Ref. Wu et al. (2021) with the measured parameters $\alpha_{\psi}$ and $\Delta\Phi_{\psi}$, the ratio of the D-wave and S-wave coupling constants $g_{D}/g_{S}$, the relative phase $\delta$ between the S-wave and D-wave, the percentage of the S-wave contribution $\Gamma_{S}/\Gamma_{\rm total}$, and the effective radius $r_{\rm eff}$ of the $\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ decay (where $\bar{L}=r_{\rm eff}\times p$, $\bar{L}$ and $p$ are the average orbital angular momentum and the relative momentum of the $\Sigma^{0}$ and $\bar{\Sigma}^{0}$ in $\psi$ C.M. system) are obtained for the first time, as shown in Table 4 together with the parameters for $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ calculated in Ref. Wu et al. (2021). What is particularly intriguing is that, as speculated in Ref. Wu et al. (2021), the $\delta$ difference between $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ is about the same as the value between $J/\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ and $\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ decays Wu et al. (2021), and both are equal to approximately $\pi$. If this is not a coincidence, the underlying mechanism should be investigated in the future. These results are crucial to understand the decay dynamics of the charmonium states and the production mechanism of the $\Sigma^{0}-\bar{\Sigma}^{0}$ pairs and provide useful information for hunting down excited nucleon resonances Wu et al. (2021).
@@ Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. The authors would like to extend thanks to Prof. Jusak Tandean, Dr. Shu-Ming Wu and Dr. Jia-Jun Wu for useful discussion and helpful advice. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406300, 2020YFA0406400; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12165022, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12342502, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; Yunnan Fundamental Research Project under Contract No. 202301AT070162; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; The Knut and Alice Wallenberg Foundation, Sweden; The Swedish Foundation for International Cooperation in Research and Higher Education (STINT); U. S. Department of Energy under Contract No. DE-FG02-05ER41374.